OpenStudy (sydthekid913):

Evaluate the following expression for x=1 and y=-3 3y^2x+x-y

OpenStudy (sydthekid913):

@haleyelizabeth2017 @sleepyjess @Bookworm14 I know this is SUPER easy and I'm sorry to bug yall but no matter what I do I cannot get the right answer :/

OpenStudy (sydthekid913):

here are my options A. -29 B. -16 C. 31 D. 25

OpenStudy (sleepyjess):

plug in 1 for x and -3 for y then use PEMDAS does it look like this? \(3y^{2x}+x-y\)

OpenStudy (sydthekid913):

no the x is not an exponent

OpenStudy (nfcfox):

Let's plug x and y in. 3(-3)^2(1) +1-(-3) Remember PEMDAS. Parentheses, exponents, multiplication, division, addition, subtraction. We can jump right into multiplication in the exponent so we get 3(-3)^2 Now we do the actual exponent. -3 squared is simply 9. Now we can multiply that to the 3 on the outside. We get 27. Our equation is now 27+1+3 (The plus 3 was just the keep change change. Add! We get 31. The answer is C. 31

OpenStudy (nfcfox):

If the x isn't an exponent then just add 1 more. We simply add one more x which is 1. The answer would be 32 if that's what you are saying but are you sure the x isn't in the exponent?

OpenStudy (sydthekid913):

\(\bbox[1pt,#FF0000]{\bbox[1pt,#FF1700]{\bbox[1pt,#FF2E00]{\bbox[1pt,#FF4500]{\bbox[1pt,#FF5C00]{\bbox[1pt,#FF7300]{\bbox[1pt,#FF8A00]{\bbox[1pt,#FFA100]{\bbox[1pt,#FFB800]{\bbox[1pt,#FFCF00]{\bbox[1pt,#FFE600]{\bbox[1pt,#FFFD00]{\bbox[1pt,#FFff00]{\bbox[1pt,#E8ff00]{\bbox[1pt,#D1ff00]{\bbox[1pt,#BAff00]{\bbox[1pt,#A3ff00]{\bbox[1pt,#8Cff00]{\bbox[1pt,#75ff00]{\bbox[1pt,#5Eff00]{\bbox[1pt,#47ff00]{\bbox[1pt,#30ff00]{\bbox[1pt,#19ff00]{\bbox[1pt,#02ff00]{\bbox[1pt,#00ff00]{\bbox[1pt,#00ff17]{\bbox[1pt,#00ff2E]{\bbox[1pt,#00ff45]{\bbox[1pt,#00ff5C]{\bbox[1pt,#00ff73]{\bbox[1pt,#00ff8A]{\bbox[1pt,#00ffA1]{\bbox[1pt,#00ffB8]{\bbox[1pt,#00ffCF]{\bbox[1pt,#00ffE6]{\bbox[1pt,#00ffFD]{\bbox[1pt,#00ffff]{\bbox[1pt,#00FDff]{\bbox[1pt,#00E6ff]{\bbox[1pt,#00CFff]{\bbox[1pt,#00B8ff]{\bbox[1pt,#00A1ff]{\bbox[1pt,#008Aff]{\bbox[1pt,#0073ff]{\bbox[1pt,#005Cff]{\bbox[1pt,#0045ff]{\bbox[1pt,#002Eff]{\bbox[1pt,#0017ff]{\bbox[1pt,#0000ff]{\bbox[1pt,#0200ff]{\bbox[1pt,#1900ff]{\bbox[1pt,#3000ff]{\bbox[1pt,#4700ff]{\bbox[1pt,#5E00ff]{\bbox[1pt,#7500ff]{\bbox[1pt,#8C00ff]{\bbox[1pt,#A300ff]{\bbox[1pt,#BA00ff]{\bbox[1pt,#D100ff]{\bbox[1pt,#E800ff]{\bbox[1pt,#FF00ff]{\bbox[1pt,#FF00FD]{\bbox[1pt,#FF00E6]{\bbox[1pt,#FF00CF]{\bbox[1pt,#FF00B8]{\bbox[1pt,#FF00A1]{\LARGE\mathbb{\color{black}{Thank you!!!\ }}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}\) Courtesy of @sleepyjess

OpenStudy (nfcfox):

What the hell seriously? He didn't even solve it for you and work it out for you, wow.

OpenStudy (sydthekid913):

*Thank you!!!

OpenStudy (sleepyjess):

(3(-3)^2)*1 +1-(-3) -3^2=9 3*9=27 27*1=27 1-(-3)=1+3=4 27+4=31

OpenStudy (sleepyjess):

what are you talking about @nfcfox

OpenStudy (sydthekid913):

I was saying thank you to you @nfcfox but sleepyjess made the LaTeX please refrain from using inappropriate language thank you:)

OpenStudy (haleyelizabeth2017):

Sorry I couldn't help @SydtheKid913, I got really busy. And @nfcfox, it wouldn't even matter if @sleepyjess solved it or not, people can still be polite and say thank you. So that was WAY uncalled for. Please try not to misunderstand something like this again because it sometimes offends people...Thank you for taking your time to read this and I'm sorry, but it's the truth. Bye @sleepyjess and @SydtheKid913.

OpenStudy (nfcfox):

@haleyelizabeth2017 no one asked for your opinion and if you don't get the slightest jealous when you think you are not getting credit for what you did then you must not care

OpenStudy (sydthekid913):

but at least you did get a medal\(\color{lime}{\Huge\ddot\smile}\)

OpenStudy (haleyelizabeth2017):

Just be quiet...please. I don't care if you didn't ask for my opinion because I was just defending my friend. And yes I do care, I just am not selfish unlike you. I'm sorry to say that but that's how you are acting right now. I'm done with this conversation. You should probably think about what I just said. I don't care for selfish people and YOU are one. \(\color{aqua}{\huge\ddot\frown}\)

OpenStudy (nfcfox):

@SydtheKid913 yes and ty for that, it was my bad for misunderstanding but I can't stand it when I don't get credit for something. @haleyelizabeth2017 I would be selfish if I wanted to steal credit from a friend, this is not the case please quiet down

OpenStudy (haleyelizabeth2017):

you are the one that needs to quiet down! You are the one that is arguing with everything I say! I'm sorry \(\color{magenta}{\huge\ddot\frown}\)

OpenStudy (nfcfox):

@haleyelizabeth2017 you were the one coming in here, i was simply defending myself, You are also the one calling me selfish.

OpenStudy (haleyelizabeth2017):

Okay. I already apologized! Can we just drop it?!

OpenStudy (sydthekid913):

Yes Please! \(\color{gold}{\Huge\ddot\smile}\)

OpenStudy (haleyelizabeth2017):

Good thing you closed this @SydtheKid913 :/

OpenStudy (sydthekid913):

ya